When x = 2 and y = 3, z = 42. Write the function that models each of the following relationships.
1. z varies inversely with x and directly with y.
2. z varies jointly with x and y
3. z varies directly with x and inversely with the square of y

Answer by stanbon(75887)

When x = 2 and y = 3, z = 42. Write the function that models each of the following relationships.
1. z varies inversely with x and directly with y.
z = k*3/2
z = (3/2)k
--------------------
2. z varies jointly with x and y
z = k*2*3
z = 6k
==============
3. z varies directly with x and inversely with the square of y
z = k*x/y^2
z = k*2/9
z = (2/9)k
==============
Cheers,
Stan H.